OpenStudy (anonymous):

ok. long problem here.1) 4x + 3le 30 x+3yle 21 xle 0 y=le ..... 3)next equation x + y le 8 5 + 5y le 20 x le 0 le z min. for c = 5x + 8y now,for 1) rewrite all the constraints in interslope form, list the maximum amount, including the x and the y value or the objective function, for 3)List the values of the objective found for each vortex......PLEASE HELP!!!!!

OpenStudy (anonymous):

hello????????????????

OpenStudy (ranga):

you can use the < and > symbols to make it clearer.

OpenStudy (anonymous):

\[4x + 3y \ge30, x + 3y \ge21, x \ge0 , y=\ge0\] rewrite the constraints in interslope form, list the maximum amount, including the x and the y value of the objective form

OpenStudy (ranga):

Okay, the slope intercept form of a line is: y = mx + b where m is the slope and b is the intercept. So we have to take each constraint and keep y on the left and take everything else to the right. I will do the first and you do the next. 4x + 3y >= 30 subtract 4x from both sides 3y >= -4x + 30 divide both sides by 3 y >= -4/3x + 10

OpenStudy (ranga):

Also, will it be possible for you to take a snap of the problem in the book and post it because the problem above is not exactly the way it appears in the book?